OpenStudy (anonymous):

solve and find domain 1/x=1/6-1/2x

jhonyy9 (jhonyy9):

the second x is on denominator or numerator ?

OpenStudy (anonymous):

denominator

jhonyy9 (jhonyy9):

so firstly x not can be zero

jhonyy9 (jhonyy9):

6=x-3 9=x

OpenStudy (anonymous):

can i see your work to compare

OpenStudy (mathmagician):

yes, i made a mistake. the actual answer is x=9

jhonyy9 (jhonyy9):

1/x = 1/6 -1/2x the denominator common will be 6x so than we get 6/(6x) = x/(6x) - 3/(6x) because those denominators are the same we can writing 6=x-3 so add to both sides 3 than will be 3+6=x so x=9

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

hey ya thanks alot

jhonyy9 (jhonyy9):

ok bye